OpenStudy (anonymous):

state the direction of the opening for the realation y=-12x^2?

OpenStudy (anonymous):

@bubblegumpop<3

OpenStudy (anonymous):

*double hugs*

OpenStudy (anonymous):

any answer choices?

OpenStudy (anonymous):

nope thats how the question was witten

OpenStudy (anonymous):

i got: x= \[\frac{ \sqrt{-3y} }{ 6 },- \frac{ \sqrt{-3y} }{ 6 }\]

OpenStudy (anonymous):

<3